Find the solution of the following couple of different differential equations x^'=4x-y and y^'=-4x+4y when t=0 ,x(0)=1 and y(0)=1


1
Expert's answer
2021-02-23T09:05:50-0500

"\begin{cases}
x^\prime =4x-y
\\
y^\prime=-4x+4y
\end{cases}" and "x(0)=y(0)=1" .


From the first equation we get "x^{\prime \prime}=4x^\prime -y^\prime =4x^\prime -\left(-4x+4y \right)=4x^\prime +4x-4y=4x^\prime +4x-4(4x-x^\prime )=4x^\prime +4x-16x+4x^\prime"


We have second order differential equation "x^{\prime \prime}-8x^\prime +12x=0" .

The characteristic equation is "\lambda ^2-8\lambda +12=0" .

The roots are "\lambda _1=2,\" "\lambda _2=6" .

The general solution is "x(t)=c_1e^{2t}+c_2e^{6t}" .

Then we can find "y=4x-x^\prime =4\left( c_1e^{2t}+c_2e^{6t} \right) -\left(2c_1e^{2t}+6c_2e^{6t}\right)= 2c_1e^{2t}-2c_2e^{6t}" .

Using "x(0)=y(0)=1" , we get a system:

"\begin{cases}
1=c_1+c_2
\\
1=2c_1-2c_2
\end{cases}" "\begin{cases}
2=2c_1+2c_2
\\
1=2c_1-2c_2
\end{cases}" "\begin{cases}
c_1=\tfrac{3}{4}
\\
c_2=\tfrac{1}{4}
\end{cases}"


Answer: "x(t)=\tfrac{1}{4}\left(3e^{2t}+e^{6t}\right), \ \ y(t)=\tfrac{1}{2}\left(3e^{2t}-e^{6t}\right)" .